# Advanced Encryption Standard AI artificial intelligence

### v1

@@ 4.2 Arithmetic Structures
For both Xilinx and Lattice FPGAs, arithmetic operations are often implemented using carry gates. The characteristics of these carry gates can, however, substantially differ between vendors. Hence, different approaches for candidate detection are required depending on the FPGA architecture.
@@
A salient advantage of this methodology is its generic nature that relieves the user from the intricate task of manually crafting exact candidates. So far, we implemented this structural approach for both Xilinx and Lattice iCE40 platforms. To integrate other FPGA architecture, the structural characteristics of that architecture must be investigated and techniques to compile structural candidates must be implemented. This can be done in a matter of hours.
@@
Having identified a set of structural candidates, we attempt to determine their implemented operations using only functional methods to remain architecture agnostic. We first determine bit order, operand memberships, and control signals for each structural candidate. Afterwards, we check against functional models from a library of known arithmetic operations using an SMT solver.
@@
We verified functionality under at least one control assignment for 96% of the carry chains across our benchmarks, see Table III. Issues only arose when carry chains where strongly interwoven with surrounding control and data-path logic, which sometimes prevented extraction of correct structural candidates. Depending on the benchmark, we automatically verified that 3% to 91% (average: 32%) of all combinational gates implement an arithmetic operation. The variation in coverage is the result of different kinds of benchmark implementations. For signal processing designs like the hilbert_transform, canny_edge_detector, and the fft64, the data path mostly comprises of arithmetic operations. central processing unit (CPU) implementations such as ibex, icicle, and simple_risc_v contain fewer carry chains and more control logic, which explains lower numbers in detected arithmetic operations and overall combinational coverage.
4.3 Data Path
The idea behind the dataflow analysis methodology of DANA [42] is, by nature, independent of the target gate type. In our case study, we leveraged DANA to identify registers and developed a separate, but related, technique to group MUX gates into word-level structures. Similar to DANA, we used control signals as well as predecessor and successor gates to identify word-level MUXes. Accordingly, to generalize our data-path analysis, we extended DANA to take on arbitrary, user-defined gate types, including MUXes, instead of refining our initial approach. In addition, we added the option to provide known word-level structures of any kind to DANA to support word-level recovery of the target gate type. For example, in our case study DANA could only search for registers and only previously identified registers could be provided as known groups. Now, DANA can also reconstruct word-level MUXes while taking previously identified registers and arithmetic operations into account.
@@ 4.3 Data Path
To evaluate the success of register recovery, we followed the evaluation proposed for DANA [42]. We generated a ground truth by leveraging symbols left in the gate-level netlists of our benchmark by the synthesizer. Then, we use normalized mutual information (NMI) and purity as metrics to compare our refined approach against the original DANA implementation. To this end, we fed known word-level structures such as BRAMs, DSPs, and arithmetic operations to DANA as additional predefined knowledge to take into account during analysis. This marginally improved NMI and purity for most benchmarks, raising the NMI by up to 0.09 the the purity by up to 0.24, see Table III.
We also employed DANA to recover word-level MUXes. However, by nature, no ground truth for MUXes exists, as they are implicitly created from HDL control flow structures by the synthesizer and then implemented in LUTs alongside other logic. Therefore, it is, at best, ambiguous which HDL constructs would be translated into a MUX and accurately measuring the quality of our results is impossible. Still, we provide an intuition for their quality by investigating the sizes of recovered MUXes. MUXes of sizes that divide the bit-size of the datapath or are multiples of a byte are likely to be correct. To this end, we plotted the distributions of the top 5 MUX sizes for all benchmarks in Appendix C.
4.4 Bit Order
Transferring our bit-order propagation algorithm from the case study to other FPGA architectures was straight-forward, because the algorithm was already generic by nature. It operates only on known pin groups of recovered structures and the netlist graph representation to determine successor and predecessor pin groups, which is applicable independent of the underlying gate types. However, CPUs netlists tended to confuse our algorithm due to the interwoven data path of ALUs. To improve bit-order propagation results, we extended our approach for consensus finding between conflicting bit orders propagated to the same pin group. To this end, we now apply three consensus finding mechanisms that try to recover a consecutive bit order.
Consider module with a group of three input pins $[i_{0},i_{1},i_{2}]$. After propagating pin indices from structures with know bit order to those with unknown bit order, each pin is annotated with multiple index candidates coming from different sources. Here, $[0,1,2]$ is a list of indices gathered from the same source structure, denoting that $i_{0}$ should receive index 00, $i_{1}$ index $1$, and $i_{2}$ index $2$. We use ???? to indicate missing index information for the respective pin and $X$ to show that the respective pin is no longer considered.
@@ 4.4 Bit Order
To evaluate our bit-order propagation, we considered the bit orders given for DSPs and BRAMs and the ones recovered for operands of arithmetic operations to be correct and started propagation from there. A reliable bit-order ground truth—given as labels left by the synthesizer—is only available for registers. For evaluation, we created registers based on the ground truth used to evaluate DANA. Furthermore, we only consider registers containing more than three FFs, as smaller ones often belong to the control path. In Table III, we report the absolute number of pin groups we consider for bit-order propagation, the share of initially ordered pin groups, and the total share of ordered pin groups after propagation. In addition, we report the proportion of ordered pin groups that are correct according to our ground truth. On average, we reconstructed 87% of bit orders across our benchmarks with 99% of them being correct, demonstrating the effectiveness of our method. In line with expectations, we observed that the quality of bit-order propagation results improved with the number of structures featuring an inherently known bit order.
4.5 Guided Symbolic Execution
@@
Manually tracing signals across clock cycles in simulation to identify the sources of DSP inputs in our case study turned out not to be scalable as it is tedious and error prone, emphasizing the need for automation. Hence, we implemented a symbolic execution (SE) approach that is guided by concrete control values obtained from virtual probing or simulation to avoid state explosion. This approach can be helpful beyond DSP analysis to semi-automatically investigate a circuit’s dynamic behavior.
@@
Our approach allows symbolically evaluating any data signal and tracing back to its origin to automatically generate equations describing that signal’s behavior over time. These equations depend only on previously defined endpoints such as global inputs, constants (e.g., from memory), or registers. Thus, guided symbolic execution produces a system-level representation that abstracts away timing behavior and instead focuses on the sequence of computations. Finally, the human reverse engineer can manually simplify and interpret the extracted equations to assign high-level meaning.
@@
Starting at the cycle of interest, our guided SE approach first constructs the Boolean function describing the combinational sub-graph in front of a net. Next, it replaces all previously identified control signals with concrete values from simulation while keeping all other variables symbolic. Each variable left in the Boolean function now corresponds to an output net of a sequential gate. If the sequential gate is not one of our endpoints, it finds the last cycle in which the gate was updated and continues back-tracing from the data inputs of the gate at that cycle. This process is repeated until the resulting Boolean function only contains variables corresponding to output nets of our endpoints.
@@
Applying guided SE is highly specific to the target implementation and the goals of the reverse engineer. Therefore, it is hard to evaluate in terms of effectiveness. However, to demonstrate its power, we applied guided SE on the Maggie netlist to extract equations describing the circuit’s algorithmic behavior during execution, after performing the same task manually. Thereby, we semi-automatically extracted the same Python script we previously assembled by hand, see Section 3.3, this time only requiring limited manual intervention. Instead of manually tracing operands back through cycles of the simulation for weeks, we recovered them symbolically using guided SE. Thereby, we reduced the effort required to perform dynamic analysis on Maggie to a matter of days.
TABLE II: Open-source benchmark resource utilization in the number of gates.
Gate Type ibex hilbert_transformer canny_edge_detector icicle simple_risc_v fft64 sha256
Xilinx Lattice Xilinx Lattice Xilinx Lattice Xilinx Lattice Xilinx Lattice Xilinx Lattice Xilinx Lattice
LUT 3503 6379 304 445 2538 3298 1525 2705 1617 2876 1722 5362 2196 3935
Carry 65 295 68 255 636 2170 65 262 76 274 212 2008 104 346
FF 1937 1963 587 569 3713 4484 1009 983 1003 1235 1988 2556 1830 2378
BRAM 0 16 0 2 24 32 270 20 6 6 3 14 0 0
DSP 1 0 0 0 0 0 0 0 0 0 4 0 0 0
Other 727 78 85 78 72 68 78 59 37 31 160 124 269 90
Total 6233 8731 1044 1349 6983 10052 2947 4029 2739 4422 4089 10064 4399 6749
TABLE III: Results of our evaluation on the open-source benchmarks described in Table II and Maggie.
addition   subtraction   counter   negation   const. mul.   comparator   unknown   total   classified   NMI (original)   NMI (our)   purity (original)   purity (our)   no. of groups   initial ordered   final ordered   correct
@@
ibex Xilinx
1
0
5
0
0
0
0
6
0.03
0.98 0.98 0.97 0.97 154 0.19 0.86 1.00
Lattice
4
0
4
0
0
0
1
9
0.04
0.96 0.97 0.97 0.97 164 0.25 0.78 0.98
icicle Xilinx
1
1
5
0
0
0
0
7
0.05
0.91 0.95 0.89 0.96 346 0.79 0.88 0.88
Lattice
2
0
7
0
0
1
0
10
0.11
0.85 0.94 0.90 0.92 110 0.30 0.57 1.00
simple_risc_v Xilinx
3
1
4
0
0
2
0
10
0.12
0.97 0.96 0.98 0.95 117 0.29 0.85 1.00
Lattice
2
1
4
0
0
6
1
14
0.15
0.98 0.98 0.98 0.96 136 0.32 0.83 1.00
canny_edge_detector Xilinx
59
0
1
5
46
3
4
118
0.82
0.81 0.81 0.61 0.64 624 0.46 0.90 0.99
Lattice
68
0
2
5
30
5
8
118
0.60
0.81 0.83 0.68 0.77 813 0.43 0.76 1.00
fft64 Xilinx
25
5
0
2
4
0
4
40
0.23
0.89 0.97 0.73 0.97 349 0.36 0.97 0.97
Lattice
93
4
7
0
4
0
1
109
0.44
0.88 0.94 0.76 0.92 622 0.57 0.87 1.00
hilbert_transformer Xilinx
7
10
0
0
0
0
0
17
0.91
0.85 0.95 0.68 0.90 126 0.43 1.00 1.00
Lattice
7
10
0
0
0
0
0
17
0.89
0.90 0.94 0.74 0.85 126 0.46 1.00 1.00
sha256 Xilinx
10
0
0
0
0
0
3
13
0.14
0.91 0.96 0.75 0.87 120 0.28 0.95 1.00
Lattice
11
0
1
0
0
0
0
12
0.10
0.94 0.97 0.81 0.90 170 0.22 0.98 1.00
maggie Lattice
3
0
25
0
0
12
2
42
0.15
N/A N/A N/A N/A 315 0.47 0.78 N/A
5 Discussion
We begin by revisiting our research questions, followed by a discussion on the threats of FPGA reverse engineering, along with challenges and prospects of potential defenses. Finally, we revisit related work and elaborate on limitations and future research directions.
5.1 Revisiting Research Questions
@@
In our case study, we successfully recovered the algorithm controlling the Taptic Engine of the iPhone 7. To answer RQ1 and provide insights into the effort required to commit IP theft, we now share our experiences.
@@
While netlist recovery was solved following a known path with suitable open-source tooling, netlist analysis presented challenges that took months to complete. In particular, the lack of hierarchy (C1), loss of data types (C2), and synthesizer optimizations (C3) initially hampered structural analysis, but could be dealt with once patterns were identified. This was achieved through an interplay of manual and automated reverse engineering by first searching for repeating structures and then systematically automating detection thereof. While we automated substantial parts of the word-level reconstruction, manual inspection and correction were inevitably necessary to interpret results and address errors.
@@
Classification of control logic (C4) had to be mostly done by hand as existing techniques only provided limited insights, see Section 3.2. In the end, we bypassed in-depth analysis of FSM state graphs and their interplay with the data path and other FSMs by relying on virtual probing. Obtaining the I/O recordings required for this purpose presented engineering challenges on its own, see Figure 7, but these are situational and depend on the PCB. The main challenge for recovering the implemented algorithm was understanding the dynamic circuit behavior (C6), which almost entirely depended on external data (C5). To this end, simulation results were manually analyzed over many weeks. Finally, understanding the implemented algorithm required domain knowledge beyond our reverse engineering expertise (C7) and required a signal processing expert.
@@
Throughout our case study, validating intermediate results within the black-box setting we were operating in presented another significant challenge. Subsequent stages always depended on previous findings, risking error propagation. This concern was particularly pronounced during word-level reconstruction and algorithmic recovery, where our efforts often felt akin to groping in the dark. Identifying and tracing errors was sometimes impossible. While the SMT-based verification of arithmetic operations offered an initial anchor point, this alone could not fully resolve the issue. In this regard, insights derived from virtual probing and the recorded I/O played a crucial role in achieving end-to-end validation of our reverse engineering result. Virtual probing was also instrumental in interpreting the data flowing through Maggie, enabling us to generate plots of processed data and extract equations representing the implemented algorithm.
@@
An overshadowing issue of our entire case study was the lack of adequate, openly available tooling to automatically investigate the recovered netlist at scale. Developing respective tooling, e.g., for netlist simulation, required immense efforts and was the main reason for many of the months spent on conducting the case study.
@@
In light of RQ2, we developed fully automated techniques that address many of the challenges we encountered in our case study. Furthermore, we demonstrated their applicability across FPGA architectures, particularly focusing on Lattice and Xilinx FPGAs. We chose Xilinx as our second evaluation platform as they are the market leader among FPGA vendors and their 7-series FPGAs are more complex than the Lattice iCE40 device family.
Generalizing netlist pre-processing, data-path analysis, and bit-order reconstruction required only minor tweaks, because the techniques used for our case study were already rather independent of the underlying architecture. Here, we merely focused on addressing issues arising from different implementations ranging from CPUs to DSP designs.
While the redesign of the functional arithmetic verification required significant effort and months to complete, adding support for different architectures now is a matter of hours and only requires extending the structural candidate identification. Furthermore, we showed that the labor-intensive dynamic extraction of the algorithm implemented on Maggie from its DSPs could be supported by our guided symbolic execution approach. This significantly reduced the time to get from word-level to an algorithmic description to at most a couple of days.
Still, we noticed differences in the quality of our algorithms’ results based on the analyzed kind of implementation. DSP designs similar to Maggie generally yield better results than the CPUs. We assume that this is because of the more structured data and control path in DSP designs. The higher number of arithmetic operations in a DSP designs compared to a CPU or cryptographic core allows us to automatically identify larger parts of the combinational data path. CPUs leverage varying bits of a register for a wide range of arithmetic operations and thus have a more interwoven data path, which sometimes required us to adapt our techniques. In particular, introducing majority voting to the bit-order propagation massively improved results for CPU implementations.
One challenge, however, remains unsolved: the identification and separation of control logic, see Section 3.2. We demonstrated that this issue can be bypassed using dynamic analysis, but this may not always be feasible, as it requires observing an operational system.
By generalizing our techniques we have demonstrated that some of the netlist reverse engineering challenges can be considered a one-time overhead, because the respective tooling only has to be developed once and can then be applied across FPGA implementations and architectures. Consequently, we contribute to significantly reducing the manual effort of FPGA netlist reverse engineering by providing open-source implementations of the techniques presented in this section, allowing for more a realistic evaluation of countermeasures.
5.2 FPGA Threats & Defenses

### v2

@@ 4.2. Arithmetic Structures
A salient advantage of this methodology is its generic nature, which relieves the user of the intricate task of manually crafting exact candidates. So far, we implemented this structural approach for both Xilinx and Lattice iCE40 platforms. To integrate other FPGA architecture, the structural characteristics of that architecture must be investigated and techniques to compile structural candidates must be implemented. This can be done in a matter of hours.
@@
Having identified structural candidates, we attempt to determine their implemented operations using functional methods to remain architecture-agnostic. We first determine bit order, operand memberships, and control signals for each structural candidate. Afterwards, we check against a library of functional models of known arithmetic operations using an SMT solver.
@@
We verified functionality under at least one control assignment for 94% of the carry chains across our benchmarks, see Table 2. Issues only arose when carry chains where strongly interwoven with surrounding control and data-path logic, which sometimes prevented extraction of correct structural candidates. Depending on the benchmark, we automatically verified that 1% to 91% (average: 32%) of all combinational gates implement an arithmetic operation. The variation in coverage is the result of different kinds of benchmark implementations. For signal processing designs like the canny_edge_detector and the fft64, the data path mostly comprises of arithmetic operations. CPU implementations such as ibex, icicle, and simple_risc_v contain fewer carry chains and more control logic, which explains lower numbers in detected arithmetic operations and overall combinational coverage.
4.3. Data Path
The idea behind the dataflow analysis methodology of DANA (Albartus et al., 2020) is, by nature, independent of the target gate type. In our case study, we leveraged DANA to identify registers and developed a separate, but related, technique to group MUX gates into word-level structures. Similar to DANA, we used control signals as well as predecessor and successor gates to identify word-level MUXes. Accordingly, to generalize our data-path analysis, we extended DANA to take on arbitrary, user-defined gate types, including MUXes, instead of refining our initial approach. In addition, we added the option to provide known word-level structures of any kind to DANA to support word-level recovery of the target gate type. For example, in our case study DANA could only search for registers and only previously identified registers could be provided as known groups. Now, DANA can also reconstruct word-level MUXes while taking previously identified registers and arithmetic operations into account.
@@ 4.3. Data Path
To evaluate the success of register recovery, we followed the evaluation proposed for DANA (Albartus et al., 2020). We generated a ground truth by leveraging symbols left in the gate-level netlists of our benchmark by the synthesizer. Then, we use normalized mutual information (NMI) and purity as metrics to compare our refined approach against the original DANA implementation. To this end, we fed known word-level structures such as BRAMs, DSPs, and arithmetic operations to DANA as additional predefined knowledge to take into account during analysis. This marginally improved NMI and purity for most benchmarks, raising the NMI by up to 0.14 the the purity by up to 0.30, see Table 2.
We also employed DANA to recover word-level MUXes. However, by nature, no ground truth for MUXes exists, as they are implicitly created from HDL control flow structures by the synthesizer and then implemented in LUTs alongside other logic. Therefore, it is, at best, ambiguous which HDL constructs would be translated into a MUX and accurately measuring the quality of our results is impossible. Still, we provide an intuition for their quality by investigating the sizes of recovered MUXes in Appendix D.
4.4. Bit Order
Transferring our bit-order propagation algorithm from the case study to other FPGA architectures was straightforward because the algorithm was already generic by nature. It operates only on known pin groups of recovered structures and the netlist graph representation to determine successor and predecessor pin groups, which is applicable independent of the underlying gate types. However, CPU netlists tend to confuse our algorithm due to the interwoven data path of ALUs. To improve bit-order propagation results, we extended our approach for consensus-finding between conflicting bit orders propagated to the same pin group. To this end, we now apply three consensus-finding mechanisms that try to recover a consecutive bit order.
Consider a module with a group of three input pins $[i_{0},i_{1},i_{2}]$. After propagating pin indices from structures with known bit orders to those with unknown bit orders, each pin is annotated with multiple index candidates coming from different sources. Here, $[0,1,2]$ is a list of indices gathered from the same source structure, denoting that $i_{0}$ should receive index 00, $i_{1}$ index $1$, and $i_{2}$ index $2$. We use ???? to indicate missing index information for the respective pin and $X$ to show that the respective pin is no longer considered.
@@ 4.4. Bit Order
To evaluate our bit-order propagation, we considered the bit orders given for DSPs and BRAMs and the ones recovered for operands of arithmetic operations to be correct and started propagation from there. A reliable bit-order ground truth—given as labels left by the synthesizer—is only available for registers. For evaluation, we created registers based on the ground truth used to evaluate DANA. Furthermore, we only consider registers containing more than three FFs, as smaller ones often belong to the control path. In Table 2, we report the absolute number of pin groups we consider for bit-order propagation, the share of initially ordered pin groups, and the total share of ordered pin groups after propagation. In addition, we report the proportion of ordered pin groups that are correct according to our ground truth. On average, we reconstructed 86% of bit orders across our benchmarks with 97% of them being correct, demonstrating the effectiveness of our method. In line with expectations, we observed that the quality of bit-order propagation results improved with the number of structures featuring an inherently known bit order.
4.5. Guided Symbolic Execution
@@
Manually tracing signals across clock cycles in simulation to identify the sources of DSP inputs in our case study turned out not to be scalable as it is tedious and error-prone, emphasizing the need for automation. Hence, we implemented a symbolic execution (SE) approach that is guided by concrete control values obtained from virtual probing or simulation to avoid state explosion. This approach can be helpful beyond DSP analysis to semi-automatically investigate a circuit’s dynamic behavior.
@@
Our approach allows symbolically evaluating any data signal and tracing it back to its origin to automatically generate equations describing that signal’s behavior over time. These equations depend only on previously defined endpoints such as global inputs, constants (e.g., from memory), or registers. Thus, guided symbolic execution produces a system-level representation that abstracts away timing behavior and instead focuses on the sequence of computations. Finally, the human reverse engineer can manually simplify and interpret the extracted equations to assign high-level meaning.
@@
Starting at the cycle of interest, our guided SE approach first constructs the Boolean function describing the combinational sub-graph in front of a net. Next, it replaces all previously identified control signals with concrete values from simulation while keeping all other variables symbolic. Each variable left in the Boolean function now corresponds to an output net of a sequential gate. If the sequential gate is not one of our endpoints, it finds the last cycle in which the gate was updated and continues back-tracing from the data inputs of the gate at that cycle. This process is repeated until the resulting Boolean function only contains variables corresponding to the output nets of our endpoints. To resolve potential recursive dependencies, our guided SE approach introduces intermediate variables to break up these dependencies when they are detected. To achieve this, we automatically identify sequential loops within the netlist and assign intermediate variables to the registers at the outputs of these loops.
@@
Applying guided SE is highly specific to the target implementation and the goals of the reverse engineer. Therefore, it is impossible to express its effectiveness in numbers. Still, we applied guided SE in a white-box case study and verified the results against a ground truth in Section 4.6 to demonstrate correctness. In a second evaluation step, we applied guided SE on the Maggie netlist to highlight its efficiency and scalability. To this end, we semi-automatically extracted a Python script describing the circuit’s algorithmic behavior during execution, after initially performing the same task manually in Section 3.3. Thereby, we reduced the required effort from multiple weeks to merely a matter of days.
4.6. White-Box Case Study
To assess the applicability of our workflow, we conducted a white-box case study on a signal-processing design similar to Maggie. This additional study is conducted because it enables a comparison between the recovered design against a known ground truth, which is impossible in the black box setting. To this end, we synthesized an open-source Hilbert transformer design (OpenCores, [n. d.]) for a Xilinx 7-series FPGA and then extracted an algorithmic abstraction from the resulting gate-level netlist using only our generalized techniques.
Word-Level Reconstruction
The initial netlist comprised 425 combinational and 619 sequential gates. After netlist pre-processing, we ended up with 1025 gates in total, 374 of them combinational and 651 sequential. Next, we recovered word-level structures using our generalized algorithms. The algorithm classified all 17 carry chains into 10 subtractions and 7 additions, identified 31 16-bit registers, and annotated 135 multi-bit pin groups to registers and arithmetic structures. When comparing against the ground truth, DANA achieved an NMI of 0.99, and our bit-order propagation assigned a correct bitorder to all pin groups. In total, we automatically assigned over 90% of all gates to word-level structures.
Algorithmic Recovery
The Hilbert transformer comes with a testbench to validate its Verilog description before implementation. The same testbench can also be used to simulate the gate-level netlist, thereby generating traces similar to those that we received from virtual probing in Section 3.3. We successfully identified all relevant control signals in the netlist by searching for nets connecting to the control pins of registers, as this is a precondition to run guided SE effectively. Next, we fed concrete values generated through simulation with the provided testbench to our guided SE algorithm for all identified control signals. Finally, our guided SE approach generated a word-level Boolean function for each global output that only depends on global inputs and the aforementioned intermediate variables introduced because of recursion.
High-Level Sensemaking
We manually translated the recovered Boolean functions into Python code, which we again handed to a domain expert who was unaware of the nature of the analyzed design. The expert identified the high-level functionality from the script and drew a block diagram that largely matched the one provided in the official documentation, see Appendix E for a comparison.
Overall, our generalized techniques enabled us to recover a high-level functional description of the analyzed design in only a few days, a task that took months to complete for the iPhone netlist. Still, significant manual effort is required despite this speedup. Overall, this white-box case study allowed us to validate the results of our workflow against a ground truth, proving its effectiveness. Furthermore, the verified success in this white-box case study supports the correctness of the analogous black-box case study, which cannot be directly verified in the same manner because the black-box setting precludes the existence of a known ground truth.
Table 2. Results of our evaluation on the open-source benchmarks described in Table 3 and Maggie.
addition subtraction counter negation const. mul. comparator unknown total classified NMI (org.) NMI (our) purity (org.) purity (our) no. of groups initial ordered final ordered correct
@@ 4.6. White-Box Case Study
ibex Lattice 5 0 3 0 0 0 1 9 0.04 0.96 0.97 0.97 0.97 165 0.25 0.39 0.91
Xilinx 3 0 0 0 0 0 3 6 0.01 0.98 0.98 0.96 0.97 148 0.16 0.86 1.00
icicle Xilinx 1 1 4 0 0 0 1 7 0.05 0.91 0.95 0.94 0.95 345 0.79 0.90 0.84
Lattice 3 0 6 0 0 1 0 10 0.12 0.86 0.94 0.90 0.94 112 0.31 0.62 0.94
simple_risc_v Xilinx 3 1 4 0 0 2 0 10 0.07 0.97 0.99 0.98 0.98 117 0.29 0.91 1.00
Lattice 2 1 4 0 0 6 1 14 0.13 0.98 0.99 0.98 0.97 136 0.32 0.95 1.00
canny_edge_detector Xilinx 57 0 1 2 49 3 6 118 0.77 0.81 0.81 0.61 0.64 619 0.46 0.91 0.99
Lattice 69 0 1 0 34 6 8 118 0.60 0.81 0.85 0.71 0.82 813 0.43 0.78 1.00
fft64 Xilinx 25 5 0 0 4 0 6 40 0.27 0.90 0.97 0.79 0.96 351 0.37 0.97 0.97
Lattice 93 4 5 0 6 0 1 109 0.40 0.88 0.94 0.77 0.91 604 0.59 0.93 0.96
sha256 Xilinx 10 0 0 0 0 0 3 13 0.26 0.91 0.96 0.75 0.87 134 0.25 0.97 1.00
Lattice 11 0 1 0 0 0 0 12 0.10 0.94 0.97 0.81 0.90 170 0.22 0.99 1.00
hilbert_transformer Xilinx 7 10 0 0 0 0 0 17 0.91 0.85 0.99 0.68 0.98 126 0.43 1.00 1.00
Lattice 7 10 0 0 0 0 0 17 0.89 0.90 0.99 0.74 0.96 126 0.46 1.00 1.00
maggie Lattice 4 0 24 0 0 15 0 43 0.16 N/A N/A N/A N/A 326 0.47 0.79 N/A
5. Discussion
5.1. Revisiting Research Questions
@@
In our case study, we successfully recovered the algorithm controlling the Taptic Engine of the iPhone 7. In response to RQ1, we now provide insights into the effort required to commit IP theft.
@@
While netlist recovery was solved following a known path with suitable open-source tooling, netlist analysis presented challenges that took months to complete. In particular, the lack of hierarchy (C1), loss of data types (C2), and synthesizer optimizations (C3) initially hampered structural analysis, but could be dealt with once patterns were identified. This was achieved through an interplay of manual and automated reverse engineering by first searching for repeating structures and then systematically automating the detection thereof. While we automated substantial parts of the word-level reconstruction, manual inspection and correction were inevitably necessary to interpret results and address errors.
@@
Classification of control logic (C4) had to be mostly done by hand as existing techniques provided limited insights, see Section 3.2. In the end, we bypassed in-depth analysis of FSM state graphs and their interplay with the data path and other FSMs by relying on virtual probing. Obtaining the I/O recordings required for this purpose presented engineering challenges on its own, see Figure 7, but these are situational and depend on the PCB. The main challenge for recovering the implemented algorithm was understanding the dynamic circuit behavior (C6), which almost entirely depended on external data (C5). To this end, simulation results were manually analyzed over many weeks. Finally, understanding the implemented algorithm required domain knowledge beyond our reverse engineering expertise (C7) and called for a signal processing expert.
@@
Throughout our case study, validating intermediate results within the black-box setting we were operating in presented another significant challenge. Subsequent stages always depended on previous findings, risking error propagation. This concern was particularly pronounced during word-level reconstruction and algorithmic recovery, where our efforts often felt akin to groping in the dark. Identifying and tracing errors was sometimes impossible. While the SMT-based verification of arithmetic operations offered an initial anchor point, this alone could not fully resolve the issue. In this regard, insights derived from virtual probing played a crucial role in achieving end-to-end validation of our reverse engineering result. Virtual probing was also instrumental in interpreting the data flowing through Maggie, enabling us to generate plots of processed data and extract equations representing the implemented algorithm.
@@
An overshadowing issue of the entire case study was the lack of adequate, openly available tooling to automatically investigate the recovered netlist at scale. Developing respective tooling, e.g., for netlist simulation, required immense efforts and was the main reason for many of the months spent on conducting the case study.
@@
In light of RQ2, we developed fully automated techniques that address many of the challenges we encountered in our case study. Furthermore, we demonstrated their applicability across FPGA architectures, particularly focusing on Lattice and Xilinx FPGAs. We chose Xilinx as our second evaluation platform as they are the market leader among FPGA vendors and their 7-series FPGAs exhibit higher complexity than the Lattice iCE40 device family.
Generalizing netlist pre-processing, data-path analysis, and bit-order reconstruction required only minor tweaks, as the techniques from our iPhone case study were already independent of the FPGA architecture. Here, we focused on addressing issues arising from different implementations ranging from CPUs to DSP designs.
While redesigning the functional arithmetic verification required significant effort and took months to complete, adding support for different architectures is now a matter of hours and only requires extending the structural candidate identification. Furthermore, we showed that the labor-intensive dynamic extraction of an algorithmic description from Maggie can be sped up by our guided SE approach. This significantly reduced the time to get from word-level to an algorithmic description to at most a couple of days.
Still, we noticed that the quality of the results of our techniques varied depending on the kind of implementation we analyzed. DSP designs similar to Maggie generally yield better results than CPUs, presumably because the data and control paths in DSP designs exhibit more structure. As DSP designs feature more arithmetic operations than, e.g., central processing units, we identify larger parts of their combinational data path. CPUs leverage varying bits of a register for a wide range of arithmetic operations and thus have a more interwoven data path, which sometimes required us to adapt our techniques. In particular, introducing majority voting to the bit-order propagation improved results for CPU implementations.
One challenge, however, remains unsolved: the identification and separation of control logic, see Section 3.2. We avoided this issue by using dynamic analysis, but it may not always be applicable, as dynamic analysis often requires observing an operational system.
By generalizing our techniques across different FPGA architectures and implementations, we demonstrated that some of the netlist reverse engineering challenges we identified can be considered a one-time overhead. The respective tooling only has to be developed once and can then be applied across a broad range of targets. Consequently, we contribute to significantly reducing the manual effort of FPGA netlist reverse engineering by providing open-source implementations of the techniques presented in this section, allowing for a more realistic evaluation of countermeasures.
5.2. FPGA Threats & Defenses